A jazz concert brought in ​$181,000 on the sale of 8,800 tickets. If the tickets sold for $15 and $25 ​each, how many of each type of ticket were​ sold?
Answer by addingup(3677)

$15: x
$25: y
x+y = 8,800 subtract y from both sides: x = 8,800-y
and
15x+25y = 181,000 substitute for x
15(8,800-y)+25y = 181,000
Distribute the 15 (Distributive Properties of Multiplication):
132,000-15y+25y = 181,000
10y = 49000
y = 4900
So, they sold 4,900 of the $25 and 8,800-4,900 = 3,900 of the $15